OpenStudy (marcelie):

help Please !!! For the following exercises, find the x- and y-intercepts of each equation 1. f (x) = − x + 2 2. g(x) = 2x + 4

OpenStudy (anonymous):

1. x-intercept: (2,0) y-intercept: (0,2) 2.x-intercept: (−2,0) y-intercept: (0,4)

OpenStudy (marcelie):

how did you get that ?

OpenStudy (lynfran):

to find the x-intercept, we let y=0 and solve for x to find the y-intercept we let x=0 and solve for y f(x)=-x+2 y=-x+2

OpenStudy (lynfran):

so y=-x+2 for x-intercept 0=-x+2 add x to both sides we get x=2 point (2,0) for y-intercept y=-(0)+2 y=0+2 y=2 point (0,2) its the same steps for the next function

OpenStudy (marcelie):

Okay ill give it a try tight now .

OpenStudy (lynfran):

ok

OpenStudy (marcelie):

Okay i got it !!! Thank you !!!